# Higashi-Hiroshima, Hiroshima 739-8526, Japan

### v1

@@ Appendix A Exact Solution of the Euclidean-Type One-Matrix Model
$R(z)=\frac{1}{2}\Bigl[z-gz^{3}-\sqrt{(z-gz^{3})^{2}+4\{g(w_{2}^{2}+zw_{1}+z^{2})-1\}}\Bigr].$ (67)
@@
$\sqrt{(z-gz^{3})^{2}+4\{g(w_{2}^{2}+z^{2})-1\}}\overset{!}{=}g(z-b)(z-c)\sqrt{(z-a)(z+a)},\qquad a\in\mathbb{R}.$ (68)
@@ Appendix B Formal Solution of the Minkowski-Type One-Matrix Model
$R(z)$ $=\frac{1}{2}\Bigl[-i(z-gz^{3})+\sqrt{-(z-gz^{3})^{2}-4i\{g(w_{2}^{2}+z^{2})-1\}}\Bigr]$ (71)
@@
Furthermore, as emphasized repeatedly, the support of the present “eigenvalue distribution” does not need to lie on the real axis, and therefore $\alpha$ is allowed to be complex. This leads to a major difference from the Euclidean model. In the Euclidean case, the requirement that $a$ be real leads to the critical value $g_{c}=\frac{1}{12}$. For $g$ exceeding this value, the single-cut assumption breaks down and the theory undergoes a phase transition (in fact, for $g>g_{c}$ the matrix integral itself is no longer well defined). Presumably reflecting this fact, the least-squares analysis in Section 4 fails to converge in the region $g>g_{c}$. This strongly suggests that the two self-consistency conditions of the present bootstrap approximation, namely that an eigenvalue distribution exists such that $w_{n}=\int d\lambda\,\lambda^{n}\rho(\lambda)$ and that the moments $w_{n}$ satisfy the loop equations, can no longer be simultaneously satisfied. In contrast, in the Minkowski model, since no reality condition is imposed on $\alpha$, the coupling $g$ appears to have no critical value. Indeed, our numerical analysis strongly suggests that the above two self-consistency conditions are compatible in all regions except for $g\sim 0$. We also expect that the instability observed around $g\sim 0$ is most likely a numerical issue originating from the choice of the cut discussed above.

### v2

@@ Appendix A Exact Solution of the Euclidean-Type One-Matrix Model
$R(z)=\frac{1}{2}\Bigl[z-gz^{3}-\sqrt{(z-gz^{3})^{2}+4\{g(w_{2}+zw_{1}+z^{2})-1\}}\Bigr].$ (67)
@@
$\sqrt{(z-gz^{3})^{2}+4\{g(w_{2}+z^{2})-1\}}\overset{!}{=}g(z-b)(z-c)\sqrt{(z-a)(z+a)},\qquad a\in\mathbb{R}.$ (68)
@@ Appendix B Formal Solution of the Minkowski-Type One-Matrix Model
$R(z)$ $=\frac{1}{2}\Bigl[-i(z-gz^{3})+\sqrt{-(z-gz^{3})^{2}-4i\{g(w_{2}+z^{2})-1\}}\Bigr]$ (71)
@@
Furthermore, as emphasized repeatedly, the support of the present “eigenvalue distribution” does not need to lie on the real axis, and therefore $\alpha$ is allowed to be complex. This leads to a major difference from the Euclidean model. In the Euclidean case, the requirement that $a$ be real leads to the critical value $g_{c}=\frac{1}{12}$. For $g$ exceeding this value, the single-cut assumption breaks down and the theory undergoes a phase transition (in fact, for $g>g_{c}$ the matrix integral itself is no longer well defined). Presumably reflecting this fact, the least-squares analysis in Section 4 fails to converge in the region $g>g_{c}$. This strongly suggests that the two self-consistency conditions of the present bootstrap approximation, namely that an eigenvalue distribution exists such that $w_{n}=\int d\lambda\,\lambda^{n}\rho(\lambda)$ and that the moments $w_{n}$ satisfy the loop equations, can no longer be simultaneously satisfied.
In the Minkowski model, by contrast, $\alpha$ is no longer constrained to be real. Therefore, at least the same mechanism of phase transition as in the Euclidean case does not seem to occur. However, one must note that the one-cut assumption that serves as the starting point of the present discussion is not supported by a clear theoretical justification as in the Euclidean case, but rather should be regarded as a working hypothesis. If this one-cut assumption were to fail, the arguments concerning the phase transition would necessarily have to be reconsidered. At present it remains an open question whether a critical value $g_{c}$ exists in the Minkowski one-matrix model. We also expect that the instability observed around $g\sim 0$ in the numerical calculation is most likely a numerical artifact arising from the particular choice of the cut described above.